# ESGenius: Benchmarking LLMs on Environmental, Social, and Governance (ESG) and Sustainability Knowledge

### v1

@@ 3.2 ESGenius-Corpus
The ESGenius-Corpus underpins every evaluation in this work, unifying authoritative frameworks, corporate questionnaires, and peer-reviewed scientific assessments that span the full breadth of environmental, social, and governance (ESG) concerns. A quantitative snapshot of the collection appears in Table 3 for IPCC climate-science assessments, Table 4 for the GRI Standards, Tables 56 for SASB’s industry-specific disclosures, Tables 78 for IFRS/ISSB guidance, Table 9 for TCFD materials, Table 10 for CDP questionnaires, and Table 11 for the UN Sustainable Development Goals. These statistics confirm (i) broad topical coverage across all three ESG pillars, (ii) deep sector-level granularity via SASB industry standards, and (iii) an expanding climate-finance focus through IFRS-ISSB, TCFD, and CDP additions.
@@
The corpus integrates sources along three tiers: (1) Core reporting standards include the Global Reporting Initiative (GRI) Standards (Initiative, 2023), Sustainability Accounting Standards Board (SASB) Standards (Board, 2023), and International Sustainability Standards Board (IFRS-ISSB) Standards (International Sustainability Standards Board, 2023); (2) Specialized reports and frameworks extend coverage with IPCC Assessment Reports (IPCC, 2023), TCFD Guidelines (Board, 2017), and CDP questionnaires (CDP Worldwide, 2023); and (3) Overarching global targets are provided by the UN Sustainable Development Goals (United Nations, 2015).
@@
The architecture is deliberately modular, allowing rapid incorporation of emerging regulations (e.g., EU CSRD, ISSB updates), sector-specific guidance (e.g., sustainable finance taxonomies), and certification frameworks. This forward-looking design safeguards the benchmark’s relevance as the ESG landscape evolves, ensuring that ESGenius remains a reliable yardstick for measuring LLM progress on sustainability understanding.
@@
The ESGenius-Corpus comprises publicly available documents and open-access materials. For proprietary standards and frameworks (e.g., GRI, SASB, CDP), we provide references to their official source locations, allowing users to access them directly, with only brief excerpts referenced for academic purposes. The corpus documentation explicitly lists included sources and provides detailed guidance on accessing external materials, ensuring reproducible research while adhering to intellectual property rights. To protect privacy, the corpus excludes all personal and sensitive information through a rigorous filtering process - sections containing individual employees, survey respondents, or contact information are systematically removed. The dataset’s anonymity is verified through both automated filters and comprehensive manual auditing.
@@ 4 Experiments and Results
Through ESGenius, we benchmark 50 LLMs spanning open-source checkpoints—DeepSeek Guo et al. (2025), Meta-Llama 3 Grattafiori et al. (2024), Google Gemma 3 Team et al. (2025), and Alibaba Qwen 2.5/3 Yang et al. (2024, 2025) from Hugging FaceHugging Face (2024) and Transformers libraryWolf et al. (2020)—as well as proprietary APIs (GPT, DeepSeek, Qwen-Max). All experiments run on a DGX A100 node (4 $\times$ 80 GB A100 GPUs) with fixed random seeds (seed=42) for reproducibility. Our evaluation suite comprises two progressively stronger settings: (1) Zero-shot prompting4.1.1), which probes the intrinsic ESG knowledge encoded during generic pre-training, and (2) Long Context RAG4.1.2), a RAG baseline that prepends relevant evidence to each prompt as long context. The RAG experiments are conducted on 43 open-source models.
For each model–setting combination, we compute exact-match accuracy on single-answer questions. Since LLMs may return raw responses that do not directly conform to our required format, we perform response validation to ensure valid multiple-choice answers (see details in Appendix A.2). This produces comprehensive leaderboards in Table 2 for both zero-shot and RAG approaches (full results in Appendix C Table 2).
@@ 4.1.1 Zero-Shot Prompting
Each model is provided the canonical question template (see Appendix B) and an example zero-shot question is shown in Appendix A.3. Response generation employs greedy decoding (temperature=0, top_p=1, top_k=0), with a maximum sequence length of 1024 tokens.
@@
Models are evaluated in half precision (FP16) using standard batched inference on GPU. Results are logged incrementally with error handling to ensure robustness.
4.1.2 Long Context RAG
@@
For each question, the pre-linked source text from ESGenius-QA is retrieved from the ESG knowledge corpus. This text serves as context, prepended to the prompt template (see Appendix B.2) and an example zero-shot question is shown in Appendix A.3. This simple but effective approach ensures that models have access to relevant domain knowledge during inference.
@@
The evaluation pipeline closely mirrors the zero-shot setup, with the key difference being the augmented prompts. We implement context injection by prepending the retrieved source text to each question prompt before feeding it to the model. The same decoding parameters (temperature=0, top_p=1) and evaluation metrics are used.
@@ 4.2 Main Results
Drawing from the experimental results presented in Table 2, we highlight three key findings and provide detailed analysis in the following sections.
@@ 4.2.1 Key Findings
(1) ESG Concept Understanding Remains Challenging for Current LLMs. Our evaluation reveals that the top-performing model, “o3”, attained an accuracy of $0.7254$, whereas the majority of other models scored approximately $0.65$. This performance level is notably comparable to, or lower than, that observed in benchmarks from other domains Guha et al. (2023), underscoring the challenges current LLMs face in accurately comprehending and reasoning about complex ESG concepts. Further details are provided in Section 4.2.2.
(2) Smaller Models Augmented with Domain-Specific RAG Can Surpass Larger Zero-Shot Models. Performance evaluation demonstrates that applying RAG significantly enhances the capabilities of smaller models on ESG question answering. For example, the accuracy of “Qwen3 4B” increased from $0.6188$ (zero-shot) to $0.7905$ using RAG. Similarly, “Qwen2.5 1.5B” rose from $0.5484$ (zero-shot) to $0.6972$. These improvements demonstrate that targeted retrieval is more effective than raw model scale for complex domain-specific knowledge.
(3) Reasoning Capabilities Enhance LLM Performance on ESG Understanding. Models augmented with explicit reasoning capabilities demonstrate enhanced performance on ESG tasks compared to non-reasoning models of comparable scale. For instance, the reasoning variant of “o3” attained an accuracy of $0.7254$, exceeding the scores of similarly sized models without such enhancements (see Figure 1). This underscores the significant benefit derived from incorporating reasoning-focused mechanisms into LLMs for specialized domains. More details in Section 4.2.4.
@@ 4.2.2 Challenging Nature of ESGenius
Our comprehensive evaluation, detailed in Table 2 (see Appendix C), reveals that the ESGenius-QA dataset is highly challenging for a broad range of contemporary LLMs. This includes proprietary APIs such as OpenAI’s GPT-4o (Zero-shot: 0.6364), DeepSeek models (“DeepSeek-R1” Zero-shot: 0.6629; “DeepSeek-V3” Zero-shot: 0.6532), and Alibaba’s Qwen2.5-Max (Zero-shot: 0.6444), as well as leading open-source families like Meta Llama (e.g., “Meta-Llama-3 8B” Zero-shot: 0.6461), Google Gemma (e.g., “Gemma-3 12B with instruction tuning” Zero-shot: 0.6514), and Alibaba Qwen (e.g., “Qwen2.5 14B” Zero-shot: 0.6734). The best zero-shot performance in the table is 0.7254 from the proprietary “o3” model, while the highest RAG-enhanced score is 0.8336 with the open-source “Gemma-3 27B with instruction tuning” model. This gap underscores both the discriminative nature of ESGenius and the benchmark’s heightened ESG reasoning demands.
@@ 4.2.3 RAG Performance
Integrating ESG-specific knowledge through long context RAG frequently yields substantial performance gains over zero-shot capabilities. For instance, the “DeepSeek-R1-Distill-Qwen” series demonstrates significant improvements across model sizes: the 1.5B model improves by 37.36% (0.3134 to 0.4305), the 7B model by 29.63% (0.5018 to 0.6505), and the 14B model by 26.07% (0.6382 to 0.8046). Other notable examples include “Gemma-3 27B” (non-instruction-tuned), which achieves a remarkable 141.57% improvement (0.2165 to 0.5229), and “QwQ 32B” with a substantial 95.23% increase (0.3900 to 0.7614). Similarly, “Qwen3 4B” improves by 27.75% (0.6188 to 0.7905), and the instruction-tuned “Qwen2.5 1.5B” model increases by 27.13% (0.5484 to 0.6972). However, RAG does not universally enhance performance. Some models actually perform worse under the RAG configuration in this study—notably, “Qwen2.5 0.5B” (standard) shows a slight decline of 1.14% (0.5458 to 0.5396), while its instruction-tuned counterpart experiences a more significant 7.48% decrease (0.5775 to 0.5343). These results indicate that RAG’s effectiveness is closely tied to model architecture and intrinsic capabilities, though it generally enables smaller models to surpass the zero-shot performance of their larger counterparts.
@@ 4.2.4 Reasoning Models Analysis
Table 2 flags “Reasoning Focus” (Rea: Yes) models. Those explicitly oriented or fine-tuned for reasoning, such as “DeepSeek-R1-Distill-Qwen” and “DeepSeek-R1-Distill-Llama” (Open Source) and proprietary offerings like “o3” (300B, Rea: Yes, ZS: 0.7254) or “o4-mini” (3B, Rea: Yes, ZS: 0.6945), frequently excel in zero-shot mode. For example, the reasoning-oriented “DeepSeek-R1-Distill-Qwen” series shows strong RAG performance and gains (14B model: 0.8046, 1.5B model improves by 37.36%), while “QwQ 32B” (Rea: Yes) also achieves a 95.23% improvement under RAG. Comparisons with non-reasoning peers reinforce this pattern. For instance, “o4 mini” (3B, Rea: Yes) reaches 0.6945 zero-shot, well above non-reasoning 3B models such as “Llama-3.2 3B” (ZS: 0.6074) or “Qwen2.5 3B” (ZS: 0.5889). These observations suggest that training or architectural choices emphasizing reasoning bolster ESG question-answering, especially for multi-step logical tasks.
@@ 4.2.5 Instruct-Tuned Models Analysis
Our analysis of instruction-tuned models (marked “I-T: Yes” in Table 2) reveals varied performance in zero-shot settings. Several models demonstrate clear benefits from instruction tuning: Google’s “Gemma-3 1B” improves substantially from 0.2421 to 0.5000, while “Gemma-3 27B” shows an even more dramatic increase from 0.2165 to 0.6356. Similar positive trends appear in Alibaba’s “Qwen2.5 0.5B” (0.5458 to 0.5775) and Meta’s “Llama-3.2 1B” (0.3609 to 0.5986). However, instruction tuning can also lead to performance degradation in some cases: “Meta Llama-3.1 8B” declines from 0.6699 to 0.6382, “Qwen2.5 1.5B” drops from 0.6391 to 0.5484, and “Qwen2.5 14B” decreases from 0.6734 to 0.6197. The impact of instruction tuning becomes particularly interesting when combined with RAG. In this context, instruction-tuned models frequently demonstrate superior improvement margins compared to their standard counterparts. For instance, the instruction-tuned version of “Gemma-3 12B” achieves 0.8380 (28.64% improvement) while its non-instruction-tuned variant reaches only 0.6857 (5.26% improvement). Similarly, the instruction-tuned “Llama-3.1 8B” attains 0.7993 (25.24% improvement) compared to the standard model’s 0.7650 (14.20% improvement). While instruction tuning may not consistently enhance zero-shot performance, its impact becomes particularly interesting when combined with RAG. In this context, instruction-tuned models frequently demonstrate superior improvement margins compared to their standard counterparts.
@@ 4.2.6 Error Analysis
Sorting the heat-map (see Section 4.2.7) columns by failure rate exposes a long tail of adversarial-grade questions that no model answered correctly. Appendix C.1 reproduces three such zero-accuracy questions. Question 432 from the SDG 2024 Report probes subtle regional disparities in per-capita renewable electricity capacity. Question 635, drawn from IPCC AR6 WG III, asks for the single most direct barrier to ESCO adoption; models consistently mis-attributed the bottleneck to regulation or awareness, missing the report’s emphasis on asymmetric information and split incentives within firms. Question 1006, based on the SASB Chemicals standard, tests the definition of a recordable workplace incident. Two broad error patterns emerge across these and similar questions: first, a tendency to rely on surface keyword overlap between prompt and distractor, and second, over-generalisation to well-known policy frameworks when a question hinges on a narrower ESG standard. These observations align with the substantial performance gains reported under RAG (§4.2), underscoring the necessity of grounding model outputs in the correct source material.

### v2

@@ 3.2 ESGenius-Corpus
The ESGenius-Corpus underpins every evaluation in this work, unifying authoritative frameworks, corporate questionnaires, and peer-reviewed scientific assessments that span the full breadth of environmental, social, and governance (ESG) concerns. A quantitative snapshot of the collection appears in Table 3 for IPCC climate-change assessments, Table 4 for the GRI Standards, Tables 5 and 6 for SASB’s standards, Tables 7 and 8 for IFRS/ISSB guidance, Table 9 for TCFD materials, Table 10 for CDP questionnaires, and Table 11 for the UN Sustainable Development Goals (SDGs). These statistics confirm (i) broad topical coverage across all three ESG pillars, (ii) deep sector-level granularity via GRI and SASB industry standards, and (iii) an expanding climate-finance focus through IFRS-ISSB, TCFD, and CDP additions.
@@
The corpus integrates sources along three tiers: (1) Core reporting standards include the Global Reporting Initiative (GRI) Standards Initiative (2023), Sustainability Accounting Standards Board (SASB) Standards Board (2023), and International Sustainability Standards Board (IFRS-ISSB) Standards International Sustainability Standards Board (2023); (2) Specialized reports and frameworks extend coverage with IPCC Assessment Reports IPCC (2023), TCFD Guidelines Board (2017), and CDP questionnaires CDP Worldwide (2023); and (3) Overarching global targets are provided by the UN SDGs United Nations (2015).
@@
The architecture of the ESGenius-Corpus is deliberately modular, enabling the rapid integration of new materials such as emerging regulations (e.g., EU CSRD, ISSB updates), sector-specific guidance (e.g., sustainable finance taxonomies), country-level policies, corporate ESG and sustainability reports, and certification frameworks (e.g., LEED for sustainable buildings). This forward-looking design ensures that the benchmark can evolve in step with the shifting ESG landscape, safeguarding its long-term relevance and maintaining ESGenius as a reliable yardstick for tracking LLM progress in sustainability understanding.
@@
The ESGenius-Corpus comprises publicly available documents and open-access materials. For proprietary standards and frameworks (e.g., GRI, SASB, CDP), we provide references to their official source locations, allowing users to access them directly, with only brief excerpts referenced for academic purposes. The corpus documentation explicitly lists included sources and provides detailed guidance on accessing external materials, ensuring reproducible research while adhering to intellectual property rights. To protect privacy, the corpus excludes all personal and sensitive information through a rigorous manual auditing process, conducted alongside the source mapping step (§3.1) and the quality control stage for dataset curation (§3.1).
@@ 4 Experiments and Results
Through ESGenius, we benchmark 50 LLMs spanning open-source checkpoints—DeepSeek Guo et al. (2025), Meta-Llama 3 Grattafiori et al. (2024), Google Gemma 3 Team et al. (2025), and Alibaba Qwen 2.5/3 Yang et al. (2024, 2025) from Hugging Face Hugging Face (2024) and Transformers library Wolf et al. (2020)—as well as proprietary APIs (GPT, DeepSeek, Qwen). All experiments run on a DGX node (4 $\times$ 80 GB A100 GPUs) with fixed random seeds (seed=42) for reproducibility. Our evaluation suite comprises two progressively stronger settings: (1) Zero-Shot Prompting4.1.1), which probes the intrinsic ESG knowledge encoded during generic pre-training, and (2) Long-Context RAG4.1.2), a RAG baseline that prepends relevant evidence to each prompt as long context. The RAG experiments are conducted on 43 open-source models.
For each model-setting combination, we compute exact-match accuracy on single-answer questions. Because LLMs often generate raw outputs that may not directly match the required format, we apply response validation to ensure valid multiple-choice answers (Appendix A.2). The results are presented as a comprehensive leaderboard (Appendix C, Table 2) for both zero-shot and RAG settings. To complement these results, we also include a ranking bar plot in Figure 3, providing a clearer visualization of LLM performance under the zero-shot setup.
@@ 4.1 Evaluation Protocols
To rigorously assess model performance on ESGenius, we design two complementary evaluation protocols. The first, Zero-Shot Prompting, measures how well models can answer ESG-related questions without external assistance, reflecting their intrinsic knowledge and reasoning ability. The second, Long-Context RAG, evaluates models in a retrieval-augmented setting by providing each question with its corresponding evidence passage from the ESGenius-Corpus. Together, these protocols enable a transparent comparison between raw model capabilities and knowledge-grounded reasoning, thereby establishing fair and reproducible baselines for future research.
@@ 4.1.1 Zero-Shot Prompting
Each model is provided the zero-shot prompt template (Appendix B.1) and an zero-shot question example is shown in Appendix A.3.1. Response generation employs greedy decoding (temperature=0, top_p=1, top_k=0), with a maximum sequence length of 1024 tokens.
@@
Models are evaluated in half precision (FP16) using standard batched inference on GPUs. Results are logged incrementally with error handling to ensure robustness.
4.1.2 Long-Context RAG
@@
For each question, the pre-linked source passage from ESGenius-QA is retrieved from the ESG knowledge corpus. This passage is prepended to the prompt template (Appendix B.2), and an illustrative example is provided in Appendix A.3.2. This simple but effective retrieval ensures that models consistently access relevant domain knowledge during inference.
@@
The evaluation pipeline mirrors the zero-shot setup, with the key difference being the augmented prompts. Context is injected by prepending the retrieved passage to each question prompt before feeding it to the model, using the same decoding parameters (temperature=0, top_p=1) and evaluation metrics. Notably, even this straightforward lexical-matching long-context RAG method yielded measurable performance gains, underscoring the benchmark’s utility. While more sophisticated retrieval-augmented generation techniques could further improve performance, our aim was to establish a transparent, fair, and reproducible RAG baseline.
@@ 4.2 Main Results
Drawing from the experimental results presented in Table 2, we highlight three key findings and provide detailed analysis from various perspectives subsequently.
@@ 4.2.1 Key Findings
(1) ESG Concept Understanding Remains Challenging for Current LLMs. Our evaluation reveals that the top-performing zero-shot model, o3, attained an accuracy of 0.7254, whereas the majority of other models scored approximately 0.65. This performance level is notably comparable to, or lower than, that observed in benchmarks from other domains Guha et al. (2023), underscoring the challenges current LLMs face in accurately comprehending and reasoning about complex ESG concepts. Further details are provided in §4.2.2.
(2) Domain-Specific RAG Allows Smaller Models to Outperform Larger Zero-Shot Counterparts. Performance evaluation demonstrates that applying RAG significantly enhances the capabilities of smaller models on ESG question answering. For example, the accuracy of Qwen3-4B increased from 0.6188 (zero-shot) to 0.7905 (RAG). Similarly, Qwen2.5-1.5B rose from 0.5484 (zero-shot) to 0.6972 (RAG). These improvements demonstrate that targeted retrieval is more effective than raw model scale for complex domain-specific knowledge. More details are described in §4.2.3.
(3) Reasoning Capabilities Enhance LLM Performance on ESG Understanding. Models augmented with explicit reasoning capabilities demonstrate enhanced performance on ESG tasks compared to non-reasoning models of comparable scale. For instance, the reasoning variant of o3 attained an accuracy of 0.7254, exceeding the scores of similarly sized models without such enhancements (Figure 2). This underscores the significant benefit derived from incorporating reasoning-focused mechanisms into LLMs for specialized domains. More details in §4.2.4.
@@ 4.2.2 Challenging Nature of ESGenius
Our comprehensive evaluation, detailed in Table 2, reveals that the ESGenius-QA dataset is highly challenging for a broad range of contemporary LLMs. This includes proprietary APIs such as GPT-4o (zero-shot: 0.6364), DeepSeek-R1 (zero-shot: 0.6629), DeepSeek-V3 (zero-shot: 0.6532), and Qwen2.5-Max (zero-shot: 0.6444), as well as leading open-source families like Meta Llama (e.g., Meta-Llama-3-8B zero-shot: 0.6461), Google Gemma (e.g., Gemma-3-12B-Instruct zero-shot:0.6514, Qwen2.5-14B zero-shot: 0.6734). The best zero-shot performance in the table is 0.7254 from the proprietary o3 model, while the highest RAG-enhanced score is 0.8336 with the open-source Gemma-3-27B-Instruct model. This gap underscores both the discriminative nature of ESGenius and the benchmark’s heightened ESG reasoning demands.
@@ 4.2.3 RAG Performance
Integrating ESG-specific knowledge through long-context RAG frequently yields substantial performance gains over zero-shot capabilities. For instance, the DeepSeek-R1-Distill-Qwen series demonstrates significant improvements across model sizes: the 1.5B model improves by 37.36% (0.3134 to 0.4305), the 7B model by 29.63% (0.5018 to 0.6505), and the 14B model by 26.07% (0.6382 to 0.8046). Other notable examples include Gemma-3-27B, which achieves a remarkable 141.57% improvement (0.2165 to 0.5229), and QwQ-32B with a substantial 95.23% increase (0.3900 to 0.7614). Similarly, Qwen3-4B improves by 27.75% (0.6188 to 0.7905), and the Qwen2.5-1.5B-Instruct model increases by 27.13% (0.5484 to 0.6972). However, RAG does not universally enhance performance. Some models actually perform worse under the RAG configuration in this study—notably, Qwen2.5-0.5B shows a slight decline of 1.14% (0.5458 to 0.5396), while its instruction-tuned counterpart experiences a more significant 7.48% decrease (0.5775 to 0.5343). These results suggest that for some smaller architectures, additional context may function as noise or exceed their ability to leverage long sequences effectively. Overall, the effectiveness of RAG appears strongly dependent on model architecture and intrinsic capacity, yet it often allows many smaller models to outperform the zero-shot performance of much larger ones.
@@ 4.2.4 Reasoning Models Analysis
Table 2 flags “Reasoning Focus” (Rea: Yes) models. Those explicitly oriented or fine-tuned for reasoning, such as DeepSeek-R1-Distill-Qwen and DeepSeek-R1-Distill-Llama (Open Source) and proprietary offerings like o3 (300B, Rea: Yes, zero-shot: 0.7254) or o4-mini (3B, Rea: Yes, zero-shot: 0.6945), frequently excel in zero-shot mode. For example, the reasoning-oriented DeepSeek-R1-Distill-Qwen series shows strong RAG performance and gains (14B model: 0.8046, 1.5B model improves by 37.36%), while QwQ-32B (Rea: Yes) also achieves a 95.23% improvement under RAG. Comparisons with non-reasoning peers reinforce this pattern. For instance, o4-mini (3B, Rea: Yes) reaches 0.6945 zero-shot, well above non-reasoning 3B models such as Llama-3.2-3B (zero-shot: 0.6074) or Qwen2.5-3B (zero-shot: 0.5889). These observations suggest that training or architectural choices emphasizing reasoning bolster ESG question-answering, especially for multi-step logical tasks.
@@ 4.2.5 Instruct-Tuned Models Analysis
Our analysis of instruction-tuned models (marked “I-T: Yes” in Table 2) reveals varied performance in zero-shot settings. Several models demonstrate clear benefits from instruction tuning: Google’s Gemma-3-1B improves substantially from 0.2421 to 0.5000, while Gemma-3-27B shows an even more dramatic increase from 0.2165 to 0.6356. Similar positive trends appear in Alibaba’s Qwen2.5-0.5B (0.5458 to 0.5775) and Meta’s Llama-3.2-1B (0.3609 to 0.5986). However, instruction tuning can also lead to performance degradation in some cases: Meta-Llama-3.1-8B declines from 0.6699 to 0.6382, Qwen2.5-1.5B drops from 0.6391 to 0.5484, and Qwen2.5-14B decreases from 0.6734 to 0.6197. The impact of instruction tuning becomes particularly interesting when combined with RAG. In this context, instruction-tuned models frequently demonstrate superior improvement margins compared to their standard counterparts. For instance, the instruction-tuned version of Gemma-3-12B achieves 0.8380 (28.64% improvement) while its non-instruction-tuned variant reaches only 0.6857 (5.26% improvement). Similarly, the instruction-tuned Llama-3.1-8B attains 0.7993 (25.24% improvement) compared to the standard model’s 0.7650 (14.20% improvement).
@@ 4.2.6 Error Analysis
Sorting the heatmap (§4.2.8) columns by failure rate exposes a long tail of adversarial-grade questions that no model answered correctly. Appendix C.1 reproduces three such zero-accuracy cases: Question 432 from the SDG 2024 Report probes subtle regional disparities in per-capita renewable electricity capacity; Question 635, drawn from IPCC AR6 WG III, asks for the most direct barrier to ESCO adoption, where models consistently misattributed the bottleneck to regulation or awareness rather than the report’s emphasis on asymmetric information and split incentives; and Question 1006, based on the SASB Chemicals standard, tests the definition of a recordable workplace incident. Across these and similar examples, several error patterns merit closer investigation: a reliance on surface keyword overlap between prompt and distractor, and a tendency to over-generalize to well-known policy frameworks when questions hinge on narrower ESG standards, suggesting that more fine-grained analyses are needed. Future work could therefore combine quantitative error clustering with qualitative case studies, cross-model comparisons, and expert-in-the-loop probing to develop a more comprehensive understanding of the root causes of such failures.
4.2.7 Sub-topic Analysis
A finer-grained sub-topic analysis provides additional insight into model strengths and weaknesses across ESG dimensions. While the current benchmark draws from seven ESG source collections to ensure broad coverage, disaggregating results by ESG pillar (Environmental, Social, Governance) and their sub-themes (§3.1) offers more actionable perspectives. In practice, Environmental spans climate change, biodiversity, pollution control, circular economy, energy transition, etc; Social covers labor rights, diversity and inclusion, community impacts, human health and safety, equity in supply chains, etc; Governance addresses board accountability, anti-corruption, transparency, reporting compliance, risk management, etc; and Cross-Cutting themes such as just transition, Indigenous knowledge, digital governance further blurring boundaries, etc. This infrastructure lays the foundation for mapping each question to its corresponding ESG pillar or sub-theme. In future versions, we will expand this perspective by (i) reporting per-theme performance scores and (ii) discussing how this lens can guide targeted dataset expansion. As part of this process, we have begun annotating each question with an E, S, G, or Cross-subtopic label. For example, Question ID 1—“According to the IPCC AR6 Synthesis Report, which statement would most likely lead to increased vulnerability of biodiversity and ecosystem services while simultaneously reducing carbon uptake, based on the relationships described?”—is labeled E (Environmental, Biodiversity). Question ID 578—“Which initiative best promotes workplace fairness and equal opportunities across different demographic groups within an organization?”—is labeled S (Social, Diversity and Inclusion). In contrast, Question ID 1127—“When an organization determines that indirect economic impacts are a material topic, which of the following best describes the conditions under which it can omit specific disclosures without violating reporting requirements, according to GRI 203: Indirect Economic Impacts 2016?”—is labeled G (Governance, Reporting Compliance). This annotation process is ongoing and will inform future analytical releases, enabling deeper performance breakdowns across ESG pillars and sub-themes.